OpenStudy (lilai3):

This is just a check! Please please please help! Will reward medal/ fan you!!! Factor the polynomial. 72p^2 -50

OpenStudy (whpalmer4):

What are you checking? I don't see your answer here...

OpenStudy (lilai3):

Oh... okay. My answer is 2(6p+5) (6p-5). Please check if it is right.

OpenStudy (whpalmer4):

It is. Good job! Here's how I would check, if I didn't already know the answer: \[2(6p+5)(6p-5) = 2(6p*6p-5*6p+5*6p-5*5) = 2(36p^2-25) = 72p^2-50\checkmark\]

OpenStudy (agent0smith):

Yes You could check by expanding it out again.

OpenStudy (lilai3):

So, I am right?

OpenStudy (whpalmer4):

what part of "it is" didn't you understand? :-)

OpenStudy (agent0smith):

Er... two posts above... read them :P

OpenStudy (lilai3):

Oh...I didn't see whpalmer's check until now...sorry.... thank you so much to both of you, though!

OpenStudy (agent0smith):

I also gave you a medal, which was a strong hint that you were correct.

OpenStudy (lilai3):

Thanks! <3

OpenStudy (whpalmer4):

or a really imaginative wrong answer :-) But yes, beating this horse into ground, you nailed it!

OpenStudy (lilai3):

What does beating this horse into ground mean? Just wondering...and I do have to learn something new everyday, right? ;)

OpenStudy (whpalmer4):

Nice to encounter someone who doesn't need factoring taught to them :-) have you heard the expression "beating this horse to death"?

OpenStudy (lilai3):

No.

OpenStudy (whpalmer4):

or beating a dead horse?

OpenStudy (lilai3):

What does it mean?

OpenStudy (lilai3):

I think I have heard of beating a dead horse.

OpenStudy (lilai3):

I may have....possibly...*

OpenStudy (whpalmer4):

well, it suggests that you've continued past the point of any need to discuss the matter, and yet are still discussing it. We've beaten this dead horse so much, we pounded it into the ground :-)

OpenStudy (lilai3):

I see. Thanks for the explanation! Can you please help me with another similar problem?

OpenStudy (whpalmer4):

Sure, why not, though I already gave you a medal for this one, so if you want another, you'll have to post it in a separate question :-)

OpenStudy (lilai3):

Oh, okay! <3